This problem is in my college algebra for calculus class under quadratic equations. The book gives the answer but not how it is worked out. Can you please help. It says to solve 3y^3-5y^2-2y=0
Answer by edjones(8007)

3y^3-5y^2-2y=0
y(3y^2-5y-2)=0
y(3y^2-6y+y-2)=0 Factoring by grouping.
y(3y(y-2)+(y-2))=0
y(3y+1)(y-2)=0
y=0, y=-1/3, y=2
.
Ed